OpenStudy (anonymous):

The equation in point-slope form of the line which passes through (-2, 3) and is parallel to 3x + 2y = 10 is A. y - 3 = -3/2 (x + 2) B. y - 2 = -3 (x - 3) C. y - 3 = 3/2 (x - 2)

OpenStudy (anonymous):

@cambrige

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I TOLD U THE TRICK DUDE

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

USE IT LOL

OpenStudy (anonymous):

what is it agian

OpenStudy (anonymous):

WHAT IS UR EQUATION

OpenStudy (anonymous):

3x + 2y = 10

OpenStudy (anonymous):

REWRITE IT BY CHANGING THE CONSTANT TO K or any alphabet

OpenStudy (anonymous):

thats it??

OpenStudy (anonymous):

rewrite and show

OpenStudy (anonymous):

3x + 2y =k?

OpenStudy (anonymous):

NOW SUBSTITUE THE POINTS IN THE EQUATION

OpenStudy (anonymous):

(-2,3)

OpenStudy (anonymous):

ok any hints on whch one?

OpenStudy (anonymous):

U WILL GET IT

OpenStudy (anonymous):

FOLLOW MY LEAD

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

DID U FIND K???

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

okey

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

DID U GET K???

OpenStudy (anonymous):

I AM NOT ANSWERING TILL U FIND K

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

cmon man......we're losing daylight

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ik ik

OpenStudy (anonymous):

3=k

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

SO

OpenStudy (anonymous):

wait what???

OpenStudy (anonymous):

I GOT 3

OpenStudy (anonymous):

I missed that..how in the world did u get k=3??

OpenStudy (anonymous):

IDK I DID

OpenStudy (anonymous):

3(-2)+2(3)=k

OpenStudy (anonymous):

is that equal to 3??

OpenStudy (anonymous):

no solution says my calc

OpenStudy (anonymous):

DONT USE CALCULATOR

OpenStudy (anonymous):

0 its 0

OpenStudy (anonymous):

YESSSSS............finally

OpenStudy (anonymous):

now wut

OpenStudy (anonymous):

NOW EXPAND OPTION A B C

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

WHICHEVER GIVES U 3x+2y=0

OpenStudy (anonymous):

DUDE EXPAND as in multiply and stuff

OpenStudy (anonymous):

dude im in a bind i kinda need the answer

OpenStudy (anonymous):

MEDALLING WONT HELP '

OpenStudy (anonymous):

DID u EXPAND A??

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so a? idk

OpenStudy (anonymous):

now can u do it..........U will never need help again I promise u dat

OpenStudy (anonymous):

ok watuv you u say

OpenStudy (anonymous):

2(y-3)=-3(x+2)

OpenStudy (anonymous):

Expand man

OpenStudy (anonymous):

what do ye get??

OpenStudy (anonymous):

i put b

OpenStudy (anonymous):

DUDE

OpenStudy (anonymous):

XPAND OPTION A ur not doin it

OpenStudy (anonymous):

2(y-3)=-3(x+2).....I can understand the frustration.....but u will emerge victorius

OpenStudy (anonymous):

cmon....last try

OpenStudy (anonymous):

sorry man i have alot more questions cant stay on 1 to long

OpenStudy (anonymous):

Doint this is THE ANSWER DUMMY

OpenStudy (anonymous):

fine........U gave up

OpenStudy (anonymous):

2y-6=-3x-6

OpenStudy (anonymous):

which gives u ??

OpenStudy (anonymous):

ya thanks anyway thogh

OpenStudy (anonymous):

DUDe WHAT IS THE ANSWER??

OpenStudy (anonymous):

A or B or C??

OpenStudy (anonymous):

idk its a 50 question test

OpenStudy (anonymous):

OK answer is A...

OpenStudy (anonymous):

gtg...........good try though